ANHARMONIC RESONANCES AMONG LOW-LYING VIBRATIONAL LEVELS OF METHYL ISO-CYANIDE (H_{3}CNC)

\underline{P. PRACNA}, J. URBAN, J. Heyrovský Institute of Physical Chemistry, v.v.i., Academy of Sciences of the Czech Republic, 18223 Prague 8, Czech Republic; Š. URBAN, J. VARGA, Department of Analytical Chemistry, Institute of Chemical Technology, 16628 Prague 6, Czech Republic; V.-M. HORNE-MAN, Department of Physical Sciences, University of Oulu, Linnanmaa, FIN-90570 Oulu, Finland.

Vibrational levels up to 1000 cm^{−1}of H_{3}C-N≡C are currently studied in FTIR spectra together with rotational transitions within these levels. This investigation comprises the low-lying excited vibrational levels of the CNC doubly degenerate bending vibration v_{8} = 1^{±1} (267.3 cm^{−1}), v_{8} = 2^{0,±2} (524.6 cm^{−1}(A), 545.3 cm^{−1}(E)), and v_{8} = 3^{±1,±3} (792.5 cm^{−1}(A1+A2), 833.9 cm^{−1}(E)), respectively, and the next higher fundamental level of the C-N valence vibration v_{4} = 1 (945 cm^{−1}). All these vibrational levels exhibit cubic and quartic anharmonic resonances localized to moderate values of the rotational quantum number K ≤ 10. Therefore the system of rovibrational levels has to be treated as a global polyad in order to describe all the available data quantitatively. The ground state constants have been improved considerably by extending the assignments to higher J/K rotational states both in the purely rotational spectra recorded in the ground vibrational level and in the ground state combination differences generated from the wavenumbers assigned in the fundamental ν_{4} band. Similarities and differences with respect to isoelectronic molecules CH_{3}CN and CH_{3}CCH are discussed.